### v1

# Asymptotically Optimal Tests for One- and Two-Sample Problems

Abstract
In this work, we revisit the one- and two-sample testing problems: binary hypothesis testing in which one or both distributions are unknown. For the one-sample test, we provide a more streamlined proof of the asymptotic optimality of Hoeffding’s likelihood ratio test, which is equivalent to the threshold test of the relative entropy between the empirical distribution and the nominal distribution. The new proof offers an intuitive interpretation and naturally extends to the two-sample test where we show that a similar form of Hoeffding’s test, namely a threshold test of the relative entropy between the two empirical distributions is also asymptotically optimal. A strong converse for the two-sample test is also obtained.
I Introduction
We consider binary hypothesis testing in which one or both distributions are unknown. Specifically, we revisit one- and two-sample testing and study asymptotically optimal schemes for both cases. The one-sample test, also known as goodness of fit test, is to determine if a given sequence of samples follows a nominal distribution. More precisely, let $X^{n}\sim Q$, the one-sample test is the following binary hypothesis test
$\begin{split}H_{0}:Q=P,\\ H_{1}:Q\neq P,\end{split}$ (1)
where $P$ is the known nominal distribution and $Q$ is unknown under $H_{1}$. The two-sample test addresses the question of whether two independent sequences follow the same unknown distributions. Given $X^{n}\sim P$ and $Y^{n}\sim Q$ with both $P$ and $Q$ unknown, the two-sample test is a binary hypothesis test in exactly the same form as (1).
The one-sample test, often referred to as the universal hypothesis testing in the statistical literature, was formulated and solved in Hoeffding’s seminal work [7]. Hoeffding showed that the so-called likelihoood ratio test achieves the optimal type II error exponent, which is the relative entropy between the nominal distribution and the unknown true distribution under $H_{1}$. Hoeffding’s test is a threshold test of the likelihood ratio between the empirical distribution and the nominal distribution. The test statistic reduces to the simple form of the relative entropy between the empirical distribution and the nominal distribution. Through the use of a properly controlled vanishing threshold, Hoeffding showed that the test is asymptotically optimal in that it achieves the optimal type II error exponent subject to a level constraint on the type I error. Hoeffding’s work inspired a large body of follow-up research, including attempts to extend the solution to continuous distributions [13, 12].
The two-sample test, also called homogeneity testing [8] or out-of-distribution sample detection [6], has found applications in various fields [2, 9]. While the two-sample test has been studied in the statistical literature, it has attracted significant attention recently in the machine learning community thanks to the discovery of the maximum mean discrepancy (MMD) test [5] that has shown promising performance across various datasets. This has inspired extensive follow-up research; and it was recently established that MMD with suitable kernels was indeed asymptotically optimal in that it achieves the optimal type II error exponent [15, 14] when subject to a level constraint on the type I error.
In this work, we revisit both the one-sample and two-sample problems for distributions with finite support. A more streamlined achievability proof of Hoeffding’s likelihood ratio test is provided that also offers a clean interpretation. The proof, as well as the test statistic itself, naturally extends to the two-sample test. More precisely, the threshold test of the relative entropy between the two empirical distributions is shown to achieve the optimal error exponent for the two-sample test, which coincides with Renyi-divergence with $\alpha=0.5$ when the two sequences are of identical length. Finally, a strong converse for the two-sample test is obtained, i.e., any test whose type II error probability decays faster than the optimal type II error exponent will have its type I error probability approaching $1$.
II Problem Formulation
II-A Notation and Preliminaries
This section collects all the notation and preliminary lemmas used in the paper.
Take $d]=\{1,2,\dots,d\}$, $\log$ denotes base 2 log, $\ln$ denotes log base $e$, and we use $\operatorname{supp}(P)$ to describe the support of a probability distribution $P$. For two probability distributions $P,Q$ , $P<<Q$ is short hand for $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$. We use $X^{n}=(X_{1},\dots,X_{n})$ to refer to a collection of $n$ samples. We say $X^{n}\sim Q$ to denote $(X_{1},\dots,X_{n})$ being independent and identically distributed samples taken from the distribution $Q$. Let $\mathcal{P}^{d}$ be the probability simplex over $d$ elements such that
$\mathcal{P}^{d}=\left\{P=(p_{1},\dots,p_{d})\in[0,1]^{d}:\sum_{i=1}^{d}p_{i}=1\right\}.$ (2)
For $P\in\mathcal{P}^{d}$, we use $p_{\min}$ to denote the minimum non-zero probability in $P$. Denote the total variation distance, Kullback-Leibler divergence (KLD) between distributions $P,Q\in\mathcal{P}^{d}$ respectively by
$\operatorname{TV}(P_{0},P_{1})=\frac{1}{2}\left\|P_{0}-P_{1}\right\|_{1}$ (3)
$D\left(P\|Q\right)=\sum_{i=1}^{d}p_{i}\log\frac{p_{i}}{q_{i}},$ (4)
when $P<<Q$, and $\infty$ otherwise. The $\alpha$-Renyi-relative divergence for $\alpha\in(0,1)$ is defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\left(\sum_{i=1}^{d}p_{i}^{\alpha}q_{i}^{1-\alpha}\right),$ (5)
when $\left|\operatorname{supp}(P)\cap\operatorname{supp}(Q)\right|\geq 1$, and $\infty$ otherwise. The Renyi-divergence is symmetric for $\alpha=\frac{1}{2}$.
For a known distribution $P\in\mathcal{P}_{d}$ and $c>0$, we denote the KLD ball around $P$ of radius $c$ by
$\mathcal{B}(P,c)$ $=\{F\in\mathcal{P}^{d}:D\left(F\|P\right)\leq c\},$ (6)
and we use $\mathcal{B}^{C}(P,c)$ to denote its complement in $\mathcal{P}^{d}$. For $X^{n}\sim P$ and $Y^{m}\sim Q$, denote the empirical distributions of these samples as $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$. Throughout this paper, we will make heavy use of the following large deviation result [4, Theorem 12.4.1]
Theorem 1 (Sanov’s Theorem).
Let $X^{n}\sim P$ for $P\in\mathcal{P}^{d}$. Let $E\subset\mathcal{P}^{d}$. Then
$\mathbb{P}[\hat{P}_{X^{n}}\in E]\leq(n+1)^{d}2^{-nD\left(F^{*}\|P\right)},$ (7)
where $P^{*}=\operatorname*{arg\,min}_{F\in E}D\left(F\|P\right)$.
For distributions with finite support, the KLD satisfies the following continuity property [1, Corollary 5.9].
Lemma 1.
Suppose $\in\mathcal{P}^{d}, A,B<<CA,B<<C$, and $\left\|A-B\right\|_{1}\leq\varepsilon$. Then
$\left|D\left(A\|C\right)-D\left(B\|C\right)\right|$ $\leq 2\log\left(\frac{1}{c_{\min}}\right)\varepsilon+\frac{2\sqrt{2}}{\ln(2)}\sqrt{\varepsilon}$ (8)
$\triangleq g(C,\varepsilon).$ (9)
Of note, their result is described for quantum states, while to the best of our knowledge this specific bound had not been shown classically. The only change made here is that for convenience we bound the binary entropy function in their expression using the inequality
$h(x)=-x\ln x-(1-x)\ln x\leq 2\sqrt{x(1-x)}.$
For $\varepsilon$ small enough, $g(C,\varepsilon)=O\left(\sqrt{\varepsilon}\right)$.
For samples $X^{n}\sim Q$, a simple binary hypothesis test is to distinguish between the two hypotheses
$\begin{split}H_{0}:Q=P_{0},\\ H_{1}:Q=P_{1},\end{split}$ (10)
where $P_{0},P_{1}$ are known distributions. Any test can be described in terms of a function $A_{n}:[d]^{n}\to[0,1]$, so that for a given $X^{n}\in[d]^{n}$ the test declares $\hat{H}_{0}$ (accepts $H_{0}$) with probability $A_{n}(X^{n})$, and declares $\hat{H}_{1}$ (rejects $H_{0}$) with probability $1-A_{n}(X^{n})$. This definition accommodates deterministic and randomized tests. For a given test $A_{n}$, denote by $\alpha(A_{n})=\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ the type I error probability and $\beta(A_{n})=\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ the type II error probability. Under the Neyman-Pearson framework, one attempts to minimize $\beta(A_{n})$ among all tests that satisfy $\alpha(A_{n})\leq\varepsilon$ for some $\varepsilon\in(0,1)$. The following is a well known result for simple hypothesis testing [4].
Theorem 2 (Chernoff-Stein Lemma).
For a given $\varepsilon\in(0,1)$ let $\beta_{n}^{*}(\varepsilon)=\inf_{A_{n}}\left\{\beta(A_{n}):\alpha(A_{n})<\varepsilon\right\}$. We have
$\lim_{n\to\infty}\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=D\left(P_{0}\|P_{1}\right).$ (11)
Stein’s lemma specifies the optimal type II error exponent for the simple hypothesis testing problem when both distributions are known. Naturally, it serves as an upper bound on the type II error exponent for the one-sample test when the distribution under the alternative hypothesis is unknown.
II-B One-Sample Problem
For the one-sample test, we are given a known nominal distribution $P$ and samples $X^{n}\sim Q$, the one-sample test is specified in (1) and illustrated in Figure˜1.
Figure 1: One-Sample Problem
Hoeffding [7] constructed the following test
$\hat{H}=\begin{cases}H_{0}&:\hat{Q}_{X^{n}}\in\mathcal{B}(P,c_{n}),\\ H_{1}&:\hat{Q}_{X^{n}}\in\mathcal{B}^{C}(P,c_{n}),\end{cases}$ (12)
where $c_{n}n^{2}\to\infty$. Hoeffding refers to this as the likelihood ratio test as the test statistic, as $D(\hat{Q}_{X^{n}}\|P)$ is equivalent to the log likelihood ratio between the empirical distribution and the nominal distribution. He showed that if $c_{n}\to 0$, this likelhood ratio test, herein denoted by $T_{\text{Hoff},n}$, has the following probabilities:
$\alpha(T_{\text{Hoff},n})$ $=2^{-nc_{n}+O(\log n)},$ (13)
$\beta(T_{\text{Hoff},n})$ $=2^{-nD\left(P\|Q\right)+O(\log n)},$ (14)
Taking $c_{n}=O\left(\frac{\log n}{n}\right)$ ensures that $\alpha(T_{\text{Hoff},n})\leq\varepsilon$ for sufficiently large $n$, and
$\liminf_{n\to\infty}\frac{1}{n}\log\beta(T_{\text{Hoff},n})\geq D\left(P\|Q\right).$ (15)
Combined with Theorem˜2, this establishes the fact that the test $T_{\text{Hoff},n}$ is asymptotically optimal with type II error exponent $D\left(P\|Q\right)$.
II-C Two-Sample Problem
The two-sample test involves samples $X^{n}\sim P$, and $Y^{n}\sim Q$, with both $P$ and $Q$ unknown. The task is to determine if both samples are generated from the same distribution, and the hypothesis test is again in exactly the same form as (1). Figure˜2 gives a visual description of this problem.
Figure 2: Two-Sample Problem
The optimal type II error exponent of the two-sample problem was first established in [15, 14] to be
$\lim_{n\to\infty}\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=\inf_{F\in\mathcal{P}^{d}}D\left(F\|P\right)+D\left(F\|Q\right).$ (16)
The achievable scheme proposed in [15, 14] relies on metrics that metrize weak convergence. An extended Sanov’s theorem, a large deviation result developed in [15, 14] involving two independent sequences, was then used to show that the threshold test of any such metric between the two empirical distributions achieves the optimal error exponent for the two-sample setting for the general case, i.e., distributions on Polish, locally compact Hausdorff space such as $R^{d}$. In contrast, by focusing on the simple case of finite support distributions, we show that direct extension of Hoeffding’s test to the two-sample setting is asymptotically optimal.
III Results
III-A One Sample Test
Hoeffding’s test [7] controls the type I error through a vanishing threshold. As the threshold decreases, i.e., the radius of the KLD ball shrinks, the acceptance region, defined by $\mathcal{B}(P,c_{n})$ shrinks around $P$. Exploiting the continuity property of KLD in Lemma˜1 for finite alphabet distributions, one can show that the distributions within the shrinking KLD ball “look similar” from the perspective of other distributions outside of the ball, i.e., for any $Q\in\mathcal{P}^{d}$ with $Q\neq P$, when $n$ is large enough, we get
$D\left(F\|Q\right)\approx D\left(P\|Q\right),\quad\forall F\in\mathcal{B}(P,c_{n}).$ (17)
This, when combined with Sanov’s theorem (Theorem˜1), shows the optimality of Hoeffding’s test.
Theorem 3.
For the test described in Equation˜12 and $c_{n}=\Theta\left(\frac{\log n}{n}\right)$, we have, for $n$ sufficiently large,
$\alpha(A_{n})\leq\varepsilon.$ (18)
Further, if $D\left(P\|Q\right)=+\infty$, then there exists $N>0$ such that for any $n\geq N$,
$\beta(A_{n})=0.$ (19)
If $D\left(P\|Q\right)<+\infty$, then instead we get
$\lim_{n\to\infty}\frac{1}{n}\log\beta(A_{n})=D\left(P\|Q\right).$ (20)
The proof can be found in Section˜IV-A.
While the optimality property was originally established in [7], the proof, which offers a simple geometric interpretation, allows one to directly contruct a test for the two-sample problem with the desired optimality.
III-B Two Sample Test
Recall that $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$ denote respectively the empirical distributions of $X^{n}$ and $Y^{n}$. The one-sample test uses the KLD ball around the nominal distribution with a vanishing radius as the acceptance region. The rate with which it vanishes is controlled in a manner to ensure type I error constraint is satisfied, i.e., the probability that the empirical distributions fall within the KLD ball should be sufficiently large. The same geometric interpretation can be used in constructing a test that works for the two-sample setting. As the sample size increases, the empirical distributions will concentrate around the true distribution, i.e., $\hat{P}_{X^{n}}$ will likely fall within the KLD ball centered at $P$ whereas $\hat{Q}_{Y^{n}}$ will likely fall within the KLD ball centered at $Q$. If $P=Q$, then the KLD ball will overlap and one can resort to Bonferroni inequality to argue that $\hat{P}_{X^{n}}$ will concentrate around the KLD ball centered at $\hat{Q}_{Y^{n}}$ and vice versa. This motivates the following test for the two-sample problem.
$\hat{H}=\begin{cases}H_{0}&:D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)\leq c_{n},\\ H_{1}&:D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)>c_{n},\end{cases}$ (21)
where $c_{n}=\Theta(n^{-1/2})$ is chosen to satisfy a type I error constraint***Our choice of $c_{n}=\Theta(n^{-1/2})$ is arbitrary, the result holds for $c_{n}=n^{-1+\gamma}$ for any $\gamma\in(0,1)$.. The following theorem states that this test achieves the optimal type II error exponent.
Theorem 4 (Achievability).
Using the test described in Equation˜21 as $A_{n}$, we have, $\forall P\in\mathcal{P}_{d}$,
$\limsup_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{n}})\in A_{n}^{C}|X^{n},Y^{n}\sim P\right]\leq\varepsilon.$ (22)
And for $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)=\emptyset$,
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in A_{n}|(X^{n},Y^{n})\sim(P,Q)\right]=0.$ (23)
Finally, for $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$, we get
$\begin{split}-\liminf_{n\to\infty}\frac{1}{n}&\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in A_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &\geq\left[D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)\right].\end{split}$ (24)
where $F^{*}=\operatorname*{arg\,min}_{F\in\mathcal{P}^{d}}D\left(F\|P\right)+D\left(F\|Q\right)$.
A proof sketch can be found in Section˜IV-B.
Notice that the test described in Equation˜21 is not symmetric in $P$ and $Q$, while the problem itself is, both in the problem formulation and in the optimal error exponent. Indeed, the result holds when replacing $D\left(\hat{P}_{X^{n}}\|\hat{Q}_{Y^{n}}\right)$ with $D\left(\hat{Q}_{Y^{n}}\|\hat{P}_{X^{n}}\right)$, and or the minimum of the two. Using the minimum test statistic yields a better type I error while achieving the same asymptotic type II exponent.
We now introduce a strong converse for the two-sample problem.
Theorem 5 (Strong Converse).
For a given test $T_{n}$, if for a particular $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$ and $P\neq Q$, we have
$\begin{split}-&\liminf_{n\to\infty}\frac{1}{n}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in T_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &>\left[D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)\right].\end{split}$ (25)
Then
$\lim_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\not\in T_{n}|X^{n},Y^{n}\sim F^{*}\right]=1.$ (26)
A sketch of the proof can be found in Section˜IV-C.
Given $P,Q$, solving for $F^{*}=(f_{k}^{*})_{k}$ yields
$f_{k}^{*}=\frac{p_{k}^{1/2}q_{k}^{1/2}}{\sum_{k=1}^{d}p_{i}^{1/2}q_{i}^{1/2}}.$ (27)
Notice that $F^{*}$ is a geometric mixture of the two distributions. It follows that, from [11, Theorem 30], the optimal type II error exponent for the two-sample test is precisely the Renyi-relative entropy, i.e.,
$D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)=2D_{1/2}(P,Q).$ (28)
Combine Theorem˜4, Theorem˜5, and (28), the test in Equation˜21 satisfies
$\begin{split}-\lim_{n\to\infty}\frac{1}{n}\log\mathbb{P}&\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in A_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &=2D_{1/2}(P,Q).\end{split}$ (29)
IV Proofs
IV-A Proof of Theorem˜3
Proof.
To get Equation˜18, we apply Theorem˜1, giving us
$\mathbb{P}\left[\hat{Q}_{X^{n}}\in\mathcal{B}^{C}(P,c_{n})|X^{n}\sim P\right]\leq(n+1)^{d}2^{-nc_{n}}.$ (30)
Taking
$c_{n}=\frac{d\log(n+1)}{n}+\frac{\log(1/\varepsilon)}{n},$ (31)
gives us the type I error constraint $\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]\leq\varepsilon$.
To see Equation˜19, we note that the condition is equivalent to $\exists k\in[d]$ such that $p_{k}>0$ but $q_{k}=0$. A type II error occurs if $D\left(\hat{Q}_{X^{n}}\|P\right)\leq c_{n}$. Combined with Pinsker’s inequality [4], we have
$c_{n}\geq\frac{1}{2\ln(2)}\left\|\hat{Q}_{X^{n}}-P\right\|_{1}^{2},$ (32)
Then for $c_{n}<\frac{p_{\min}^{2}}{2\ln(2)}$ we must have $\hat{Q}_{X^{n}}[k]>0$, however since $Q[k]=0$ this is a probability 0 event.
From Theorem˜2, the $D(P\|Q)$ is the optimal type II error exponent for the simple hypothesis test and is thus an upper bound for the one-sample test as $Q$ is unknown. To show Equation˜20, we need only give an achievability proof. Furthermore, by Theorem˜1 it is sufficient to show that for any $\delta>0$ there is $N_{\delta}>0$ such that
$\min_{\hat{Q}_{X^{n}}\in\mathcal{B}(P,c_{n})}D\left(\hat{Q}_{X^{n}}\|Q\right)\geq D\left(P\|Q\right)-\delta,\quad\forall n\geq N_{\delta}.$ (33)
Now, note that for any $F<<Q$ with $F\in\mathcal{B}(P,c_{n})$, we have
$c_{n}\geq D\left(F\|P\right)$ $\geq\frac{1}{2\ln 2}\left\|F-P\right\|_{1}^{2}$ (34)
$\implies\left\|F-P\right\|_{1}$ $\leq\sqrt{c_{n}2\ln 2}.$ (35)
Invoking Lemma˜1, we have
$\left|D\left(F\|Q\right)-D\left(P\|Q\right)\right|\leq g\left(Q,\sqrt{c_{n}2\ln 2}\right).$ (36)
However, since $g(Q,x)$ is continuous in $x$ and $g(Q,0)=0$ we have $\exists U_{\delta}>0$ such that $\forall\tau\in[0,U_{\delta}]$, $g(Q,\tau)\leq\delta$. But since $c_{n}\searrow 0$, $\exists N_{\delta}>0$ such that
$\sqrt{c_{n}2\ln 2}\in[0,U_{\delta}],\quad\forall n\geq N_{\delta}.$ (37)
Therefore, we have
$\left|D\left(F\|Q\right)-D\left(P\|Q\right)\right|\leq\delta,$ (38)
and since this holds for all $F\in\mathcal{B}(P,c_{n})$, we have Equation˜33. ∎
IV-B Proof of Theorem˜4
Proof.
Equation˜23 is obvious from the fact that $c_{n}<\infty$.
Now we analyze Equation˜22. Let
$\delta_{n}$ $=\frac{1+\log(1/\varepsilon)}{n}+\frac{d\log n}{n}=\Theta\left(\frac{\log n}{n}\right),$ (39)
$c_{n}$ $=\delta_{n}+n^{-1/2}=\Theta(n^{-1/2}).$ (40)
Now,
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ (41)
$=\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)\geq c_{n}|\hat{P}_{X^{n}}\in\mathcal{B}(P,\delta_{n})\right]\mathbb{P}\left[\hat{P}_{X^{n}}\in\mathcal{B}(P,\delta_{n})\right]$ (42)
$+\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)\geq c_{n}|\hat{P}_{X^{n}}\in\mathcal{B}^{C}(P,\delta_{n})\right]\mathbb{P}\left[\hat{P}_{X^{n}}\in\mathcal{B}^{C}(P,\delta_{n})\right]$ (43)
$\leq\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)\geq c_{n}|\hat{P}_{X^{n}}\in\mathcal{B}(P,\delta_{n})\right]+(n+1)^{d}2^{-n\delta_{n}}$ (44)
$\leq\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)\geq c_{n}|\hat{P}_{X^{n}},\hat{P}_{Y^{n}}\in\mathcal{B}(P,\delta_{n})\right]+2(n+1)^{d}2^{-n\delta_{n}}.$ (45)
Note that $2(n+1)^{d}2^{-n\delta_{n}}\leq\varepsilon$ for $n$ sufficiently large.
Assume that $D\left(\hat{P}_{X^{n}}\|P\right),D\left(\hat{P}_{Y^{n}}\|P\right)\leq\delta_{n}$. Therefore, $\exists N_{1}>0$ such that for all $n\geq N_{1}$ we have
$\left\|\hat{P}_{Y^{n}}-P\right\|_{1}$ $\leq\frac{p_{\min}}{2}$ (46)
$\implies\hat{p}_{Y^{n},i}$ $\geq\frac{p_{\min}}{2},\quad\forall i\in[d].$ (47)
Now, using the fact that $\ln(x)\leq x-1$ for $x>0$, we get
$D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)$ $=\sum_{i=1}^{d}\hat{p}_{X^{n},i}\log\frac{\hat{p}_{X^{n},i}}{\hat{p}_{Y^{n},i}}$ (48)
$\leq\frac{1}{\ln 2}\sum_{i=1}^{d}\hat{p}_{X^{n},i}\frac{\hat{p}_{X^{n},i}-\hat{p}_{Y^{n},i}}{\hat{p}_{Y^{n},i}}$ (49)
$=\frac{1}{\ln 2}\sum_{i=1}^{d}\frac{(\hat{p}_{X^{n},i}-\hat{p}_{Y^{n},i})^{2}}{\hat{p}_{Y^{n},i}}$ (50)
$\leq\frac{2}{p_{\min}\ln 2}\left\|\hat{P}_{X^{n}}-\hat{P}_{Y^{n}}\right\|_{1}^{2}$ (51)
$\leq\frac{16}{p_{\min}}\delta_{n},$ (52)
where the last inequality comes from applying Pinsker’s inequality to our assumption on $\hat{P}_{X^{n}},\hat{P}_{Y^{n}}$. Since $c_{n}=\Theta(n^{-1/2})$ while $\delta_{n}=\Theta(\frac{\log n}{n})$, $\exists N_{2}>0$ such that $\forall n\geq N_{2}$ we have
$\frac{16}{p_{\min}}\delta_{n}<c_{n}$ (53)
Consequently, for $n\geq\max\{N_{1},N_{2}\}$ we get
$D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)\leq\frac{16}{p_{\min}}\delta_{n}<c_{n}$ (54)
$\implies$ $\mathbb{P}\left[\left.D\left(\hat{P}_{X^{n}}\|\hat{P}_{Y^{n}}\right)>c_{n}\right|\hat{P}_{X^{n}},\hat{P}_{Y^{n}}\in\mathcal{B}(P,\delta_{n})\right]=0,$ (55)
which gives us the type I error bound.
Finally, to prove Equation˜24 we use a similar argument to the proof of Equation˜20. Specifically, in order to have a type II error occur for this test, we must have $\hat{P}_{X^{n}},\hat{Q}_{Y^{n}}$ close together. And for any $\eta>0$ if they are sufficiently close, then by Lemma˜1 they will have
$\left|D\left(\hat{P}_{X^{n}}\|P\right)-D\left(\hat{Q}_{Y^{n}}\|P\right)\right|\leq g(P,\eta),$ (56)
so that
$D\left(\hat{P}_{X^{n}}\|P\right)+D\left(\hat{Q}_{Y^{n}}\|Q\right)$ (57)
$\leq D\left(\hat{Q}_{Y^{n}}\|P\right)+D\left(\hat{Q}_{Y^{n}}\|Q\right)+g(P,\eta)$ (58)
$\leq D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)+g(P,\eta).$ (59)
And since $g(P,\eta)\to 0$ as $\eta\to 0$, and $\eta$ is arbitrary we recover the result.
The existence and uniqueness of $F^{*}$ comes from the fact that we are minimizing a convex function over $\mathcal{P}^{d}$, which is a compact convex set [3]. ∎
IV-C Proof of Theorem˜5
Proof.
By the theorem statement we can assume that there is some $\varepsilon>0$ such that
$\begin{split}-&\liminf_{n\to\infty}\frac{1}{n}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{n}})\in T_{n}|(X^{n},Y^{n})\sim(P,Q)\right]\\ &=\left[D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)+\varepsilon\right].\end{split}$ (60)
For this $\varepsilon$, we can find $c_{\varepsilon}>0$ with the property that $\forall A,B\in\mathcal{B}(F^{*},c_{\varepsilon})$
$D\left(A\|P\right)+D\left(B\|Q\right)<D\left(F^{*}\|P\right)+D\left(F^{*}\|Q\right)+\varepsilon.$ (61)
The acceptance region $T_{n}$ will eventually have no overlap with $\mathcal{B}(F^{*},c_{\varepsilon})\times\mathcal{B}(F^{*},c_{\varepsilon})$, since this would violate the assumption on our type II error exponent.
Now, suppose that $X^{n},Y^{n}\sim F^{*}$, the empirical distributions must converge to the true distribution [10], i.e.
$\hat{F}_{X^{n}},\hat{F}_{Y^{n}}\overset{\mathbb{P}}{\to}F^{*}.$ (62)
Therefore we have
$\lim_{n\to\infty}\mathbb{P}[(\hat{F}_{X^{n}},\hat{F}_{Y^{n}})\in\mathcal{B}(F^{*},c_{\varepsilon})\times\mathcal{B}(F^{*},c_{\varepsilon})]=1,$
which bounds the type I error.
V Conclusion
This paper provided a more streamlined proof of the asymptotic optimality of Hoeffding’s likelihood ratio test for the one-sample problem. The test statistic and the proof itself naturally extend to the two-sample test, and we show that a simple test comparing the relative entropy between the two empirical distributions is asymptotically optimal for the two-sample problem. A strong converse was established for the two-sample test: any test whose type II error probability vanishes faster than the optimal type II error exponent necessarily has its type I error approaching one.