### v2

# Asymptotically Optimal Tests for One- and Two-Sample Problems

Abstract
In this work, we revisit the one- and two-sample testing problems: binary hypothesis testing in which one or both distributions are unknown. For the one-sample test, we provide a more streamlined proof of the asymptotic optimality of Hoeffding’s likelihood ratio test, which is equivalent to the threshold test of the relative entropy between the empirical distribution and the nominal distribution. The new proof offers an intuitive interpretation and naturally extends to the two-sample test where we show that a similar form of Hoeffding’s test, namely a threshold test of the relative entropy between the two empirical distributions is also asymptotically optimal. A strong converse for the two-sample test is also obtained.
I Introduction
We consider binary hypothesis testing in which one or both distributions are unknown. Specifically, we revisit one- and two-sample testing and study asymptotically optimal schemes for both cases. The one-sample test, also known as goodness of fit test, is to determine if a given sequence of samples follows a nominal distribution. More precisely, let $X^{n}\sim Q$, the one-sample test is the following binary hypothesis test
$\begin{split}H_{0}:Q=P,\\ H_{1}:Q\neq P,\end{split}$ (1)
where $P$ is the known nominal distribution and $Q$ is unknown under $H_{1}$. The two-sample test addresses the question of whether two independent sequences follow the same unknown distributions. Given $X^{n}\sim P$ and $Y^{n}\sim Q$ with both $P$ and $Q$ unknown, the two-sample test is a binary hypothesis test in exactly the same form as (1).
The one-sample test, often referred to as the universal hypothesis testing in the statistical literature, was formulated and solved in Hoeffding’s seminal work [9]. Hoeffding showed that the so-called likelihood ratio test achieves the optimal type II error exponent, which is the relative entropy between the nominal distribution and the unknown true distribution under $H_{1}$. Hoeffding’s test is a threshold test of the likelihood ratio between the empirical distribution and the nominal distribution. The test statistic reduces to the simple form of the relative entropy between the empirical distribution and the nominal distribution. Through the use of a properly controlled vanishing threshold, Hoeffding showed that the test is asymptotically optimal in that it achieves the optimal type II error exponent subject to a level constraint on the type I error. Hoeffding’s work inspired a large body of follow-up research, including attempts to extend the solution to continuous distributions [18, 17]. A second order analysis, which describes finite sample behavior of the error, was carried out for one-sample tests [6].
The two-sample test, also called homogeneity testing [10] or out-of-distribution sample detection [8], has found applications in various fields [2, 12]. While the two-sample test has been studied in the statistical literature, it has attracted significant attention recently in the machine learning community thanks to the discovery of the maximum mean discrepancy (MMD) test [4] that has shown promising performance across various datasets. This has inspired extensive follow-up research; and it was recently established that MMD with suitable kernels was indeed asymptotically optimal in that it achieves the optimal type II error exponent [21, 20] when subject to a level constraint on the type I error. A closely related formulation is the Gutman classification problem[5], where training sequences $X^{n}\sim P_{1}$ and $Y^{n}\sim P_{2}$ are given, along with a test sequence $Z^{m}\sim Q$. The goal is to determine if $Q=P_{1}$ or $Q=P_{2}$, using only the empirical samples. In [19] a second order asymptotic analysis was performed for the Gutman classication problem. At the same time as this work was released, a second order analysis of two-sample hypothesis testing was executed in [7].
In this work, we revisit both the one-sample and two-sample problems for distributions with finite support. A more streamlined achievability proof of Hoeffding’s likelihood ratio test is provided that also offers a clean interpretation. The proof, as well as the test statistic itself, naturally extends to the two-sample test. More precisely, the threshold test of the relative entropy between the two empirical distributions is shown to achieve the optimal error exponent for the two-sample test, which coincides with Renyi-divergence with $\alpha=\frac{1}{2}$ when the two sequences are of identical length. Finally, a strong converse for the two-sample test is obtained, i.e., any test whose type II error probability decays faster than the optimal type II error exponent will have its type I error probability approaching $1$.
II Problem Formulation
II-A Notation and Preliminaries
This section collects all the notation and preliminary lemmas used in the paper.
Take $d]=\{1,2,\dots,d\}$, $\log$ denotes base 2 logarithm, $\ln$ denotes logarithm base $e$, and we use $\operatorname{supp}(P)$ to describe the support of a probability distribution $P$. For two probability distributions $P,Q$ , $P<<Q$ is short hand for $\operatorname{supp}(P)\subseteq\operatorname{supp}(Q)$. We use $X^{n}=(X_{1},\dots,X_{n})$ to refer to a collection of $n$ samples. We say $X^{n}\sim Q$ to denote $(X_{1},\dots,X_{n})$ being independent and identically distributed samples taken from the distribution $Q$. Let $\mathcal{P}^{d}$ be the probability simplex over $d$ elements such that
$\mathcal{P}^{d}=\left\{P=(p_{1},\dots,p_{d})\in[0,1]^{d}:\sum_{i=1}^{d}p_{i}=1\right\}.$ (2)
For $P\in\mathcal{P}^{d}$, we use $p_{\min}$ to denote the minimum non-zero probability in $P$. Denote the total variation distance between distributions $P,Q\in\mathcal{P}^{d}$ by
$\operatorname{TV}(P,Q)=\frac{1}{2}\left\|P-Q\right\|_{1},$ (3)
and the Kullback-Leibler divergence (KLD)
$D\left(P\middle\|Q\right)=\sum_{i=1}^{d}p_{i}\log\frac{p_{i}}{q_{i}},$ (4)
when $P<<Q$, and $\infty$ otherwise. The $\alpha$-Renyi-relative divergence for $\alpha\in(0,1)$ is defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\left(\sum_{i=1}^{d}p_{i}^{\alpha}q_{i}^{1-\alpha}\right),$ (5)
when $\left|\operatorname{supp}(P)\cap\operatorname{supp}(Q)\right|\geq 1$, and $\infty$ otherwise. The Renyi-divergence is symmetric for $\alpha=\frac{1}{2}$.
For a known distribution $P\in\mathcal{P}_{d}$ and $c>0$, we denote the KLD ball around $P$ of radius $c$ by
$\mathcal{B}(P,c)$ $=\{F\in\mathcal{P}^{d}:D\left(F\middle\|P\right)\leq c\},$ (6)
and we use $\mathcal{B}^{C}(P,c)$ to denote its complement in $\mathcal{P}^{d}$. For $X^{n}\sim P$ and $Y^{m}\sim Q$, denote the empirical distributions of these samples as $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{n}}$. Throughout this paper, we will make heavy use of the following large deviation result [3, Theorem 12.4.1]
Theorem 1 (Sanov’s Theorem).
Let $X^{n}\sim P$ for $P\in\mathcal{P}^{d}$. Let $E\subset\mathcal{P}^{d}$. Then
$\mathbb{P}[\hat{P}_{X^{n}}\in E]\leq(n+1)^{d}2^{-nD\left(F^{*}\middle\|P\right)},$ (7)
where $F^{*}=\operatorname*{arg\,min}_{F\in E}D\left(F\middle\|P\right)$.
For distributions with finite support, the KLD satisfies the following continuity property [1, Corollary 5.9].
Lemma 1.
Suppose $\in\mathcal{P}^{d}, A,B<<CA,B<<C$, and $\left\|A-B\right\|_{1}\leq\varepsilon$. Then
$\left|D\left(A\middle\|C\right)-D\left(B\middle\|C\right)\right|$ $\leq 2\log\left(\frac{1}{c_{\min}}\right)\varepsilon+\frac{2\sqrt{2}}{\ln(2)}\sqrt{\varepsilon}$ (8)
$\triangleq g(C,\varepsilon),$ (9)
where $c_{\min}$ is the minimum non-zero probability in $C$.
Of note, their result is described for quantum states, while to the best of our knowledge this specific bound had not been shown classically. The only change made here is that for convenience we bound the binary entropy function in their expression using the inequality
$h(x)=-x\ln x-(1-x)\ln x\leq 2\sqrt{x(1-x)}.$
For $\varepsilon$ small enough, $g(C,\varepsilon)=O\left(\sqrt{\varepsilon}\right)$.
For samples $X^{n}\sim Q$, a simple binary hypothesis test is to distinguish between the two hypotheses
$\begin{split}H_{0}:Q=P_{0},\\ H_{1}:Q=P_{1},\end{split}$ (10)
where $P_{0},P_{1}$ are known distributions. Any test can be described in terms of a function $A_{n}:[d]^{n}\to[0,1]$, so that for a given $X^{n}\in[d]^{n}$ the test declares $\hat{H}_{0}$ (accepts $H_{0}$) with probability $A_{n}(X^{n})$, and declares $\hat{H}_{1}$ (rejects $H_{0}$) with probability $1-A_{n}(X^{n})$. This definition accommodates deterministic and randomized tests. For a given test $A_{n}$, denote by $\alpha(A_{n})=\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ the type I error probability and $\beta(A_{n})=\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ the type II error probability. Under the Neyman-Pearson framework, one attempts to minimize $\beta(A_{n})$ among all tests that satisfy $\alpha(A_{n})\leq\varepsilon$ for some $\varepsilon\in(0,1)$. The following is a well known result for simple hypothesis testing [3].
Theorem 2 (Chernoff-Stein Lemma).
For a given $\varepsilon\in(0,1)$ let $\beta_{n}^{*}(\varepsilon)=\inf_{A_{n}}\left\{\beta(A_{n}):\alpha(A_{n})<\varepsilon\right\}$. We have
$\lim_{n\to\infty}-\frac{1}{n}\log\beta^{*}_{n}(\varepsilon)=D\left(P_{0}\middle\|P_{1}\right).$ (11)
Stein’s lemma specifies the optimal type II error exponent for the simple hypothesis testing problem when both distributions are known. Naturally, it serves as an upper bound on the type II error exponent for the one-sample test when the distribution under the alternative hypothesis is unknown.
II-B One-Sample Problem
For the one-sample test, we are given a known nominal distribution $P$ and samples $X^{n}\sim Q$, the one-sample test is specified in (1) and illustrated in Figure˜1.
Figure 1: In the one-sample problem we know the nominal distribution $P$, and all information about $Q$ is taken from the sample $X^{n}$
Hoeffding [9] constructed the following test
$\hat{H}=\begin{cases}H_{0}&:\hat{Q}_{X^{n}}\in\mathcal{B}(P,c_{n}),\\ H_{1}&:\hat{Q}_{X^{n}}\in\mathcal{B}^{C}(P,c_{n}),\end{cases}$ (12)
Hoeffding refers to this as the likelihood ratio test, as the test statistic, $D(\hat{Q}_{X^{n}}\|P)$, is equivalent to the log likelihood ratio between the empirical distribution and the nominal distribution. He showed that if $c_{n}\to 0$, this likelihood ratio test, herein denoted by $T_{\text{Hoff},n}$, has the following probabilities:
$\alpha(T_{\text{Hoff},n})$ $=2^{-nc_{n}+O(\log n)},$ (13)
$\beta(T_{\text{Hoff},n})$ $=2^{-nD\left(P\middle\|Q\right)+O(\log n)},$ (14)
Taking $c_{n}=O\left(\frac{\log n}{n}\right)$ ensures that $\alpha(T_{\text{Hoff},n})\leq\varepsilon$ for sufficiently large $n$, and
$\liminf_{n\to\infty}-\frac{1}{n}\log\beta(T_{\text{Hoff},n})\geq D\left(P\middle\|Q\right).$ (15)
Combined with Theorem˜2, this establishes the fact that the test $T_{\text{Hoff},n}$ is asymptotically optimal with type II error exponent $D\left(P\middle\|Q\right)$.
II-C Two-Sample Problem
In the two-sample problem we are given samples $X^{n}\sim P$ and $Y^{m}\sim Q$, with both $P$ and $Q$ unknown. The task is to determine if both samples are generated from the same distribution, and the hypothesis test is in exactly the same form as (1). Here we allow $n\neq m$, but under we assume that $\lim_{n,m\to\infty}\frac{n}{n+m}=w\in\left[\frac{1}{2},1\right]$, representing potentially different sample rates for the two distributions. We make the assumption that $n>m$ without loss of generality, as we can simply exchange $X^{n}$ and $Y^{m}$ in our definition. Figure˜2 gives a visual description of this problem.
Figure 2: In the two-sample problem we have no knowledge of either true distribution, with all information about $P$ and $Q$ coming from the two samples $X^{n}$ and $Y^{m}$
For a given $n,m$, the type II error rate is
$\beta_{n,m}^{*}(\varepsilon)=\inf_{A_{n,m}}\{\beta(A_{n,m}):\alpha(A_{n,m})\leq\varepsilon\}$ (16)
The optimal type II error exponent of the two-sample problem was first established in [21, 20] to be
$\begin{split}\lim_{\begin{subarray}{c}n,m\to\infty\end{subarray}}\frac{-1}{n+m}&\log\beta_{n,m}^{*}(\varepsilon)\\ =\inf_{F\in\mathcal{P}^{d}}&wD\left(F\middle\|P\right)+(1-w)D\left(F\middle\|Q\right).\end{split}$ (17)
The achievable scheme proposed in [21, 20] relies on metrics that metrize weak convergence. An extended Sanov’s theorem, a large deviation result developed in [21, 20] involving two independent sequences, was then used to show that the threshold test of any such metric between the two empirical distributions achieves the optimal error exponent for the two-sample setting for the general case, i.e., distributions on Polish, locally compact Hausdorff space such as $R^{d}$. In contrast, by focusing on the simple case of finite support distributions, we show that direct extension of Hoeffding’s test to the two-sample setting is asymptotically optimal.
III Results
III-A One Sample Test
Hoeffding’s test [9] controls the type I error through a vanishing threshold. As the threshold decreases, i.e., the radius of the KLD ball shrinks, the acceptance region, defined by $\mathcal{B}(P,c_{n})$ shrinks around $P$. Exploiting the continuity property of KLD in Lemma˜1 for finite alphabet distributions, one can show that the distributions within the shrinking KLD ball “look similar” from the perspective of other distributions outside of the ball, i.e., for any $Q\in\mathcal{P}^{d}$ with $Q\neq P$, when $n$ is large enough, we get
$D\left(F\middle\|Q\right)\approx D\left(P\middle\|Q\right),\quad\forall F\in\mathcal{B}(P,c_{n}).$ (18)
This, when combined with Sanov’s theorem (Theorem˜1), shows the optimality of Hoeffding’s test.
Theorem 3.
For the test described in Equation˜12 and $c_{n}=\Theta\left(\frac{\log n}{n}\right)$, we have, for $n$ sufficiently large,
$\alpha(A_{n})\leq\varepsilon.$ (19)
Further, if $D\left(P\middle\|Q\right)=+\infty$, then there exists $N>0$ such that for any $n\geq N$,
$\beta(A_{n})=0.$ (20)
If $D\left(P\middle\|Q\right)<+\infty$, then instead we get
$\lim_{n\to\infty}-\frac{1}{n}\log\beta(A_{n})=D\left(P\middle\|Q\right).$ (21)
The proof can be found in Appendix˜B .
While the optimality property was originally established in [9], our proof, which offers a simple geometric interpretation, allows one to directly construct a test for the two-sample problem with the desired optimality.
III-B Two Sample Test
Recall that $\hat{P}_{X^{n}}$ and $\hat{Q}_{Y^{m}}$ denote respectively the empirical distributions of $X^{n}$ and $Y^{m}$. The one-sample test uses the KLD ball around the nominal distribution with a vanishing radius as the acceptance region. The rate with which it vanishes is controlled in a manner to ensure type I error constraint is satisfied, i.e., the probability that the empirical distributions fall within the KLD ball should be sufficiently large. The same geometric interpretation can be used in constructing a test that works for the two-sample setting. As the sample size increases, the empirical distributions will concentrate around the true distribution, i.e., $\hat{P}_{X^{n}}$ will likely fall within the KLD ball centered at $P$ whereas $\hat{Q}_{Y^{m}}$ will likely fall within the KLD ball centered at $Q$. If $P=Q$, then the KLD ball will overlap and one can resort to Bonferroni inequality to argue that $\hat{P}_{X^{n}}$ will concentrate around the KLD ball centered at $\hat{Q}_{Y^{m}}$ and vice versa. This motivates the following test for the two-sample problem.
$\hat{H}=\begin{cases}H_{0}&:D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)\leq c_{m},\\ H_{1}&:D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)>c_{m},\end{cases}$ (22)
where $c_{m}=\Theta(m^{-1/2})$ is chosen to satisfy a type I error constraint***Our choice of $c_{m}=\Theta(m^{-1/2})$ is arbitrary, the result holds for $c_{m}=m^{-1+\gamma}$ for any $\gamma\in(0,1)$.. The following theorem states that this test achieves the optimal type II error exponent.
Theorem 4 (Achievability).
Given $\varepsilon\in(0,1)$ the constraint on the type I error, using the test described in Equation˜22 as $A_{n,m}$, we have, $\forall P\in\mathcal{P}_{d}$,
$\limsup_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in A_{n,m}^{C}|X^{n},Y^{m}\sim P\right]\leq\varepsilon.$ (23)
And for $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)=\emptyset$,
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in A_{n,m}|(X^{n},Y^{m})\sim(P,Q)\right]=0.$ (24)
Finally, for $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$, we get
$\begin{split}\liminf_{n,m\to\infty}-\frac{1}{n+m}&\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in A_{n.m}|(X^{n},Y^{m})\sim(P,Q)\right]\\ &\geq\left[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)\right].\end{split}$ (25)
where $F^{*}=\operatorname*{arg\,min}_{F\in\mathcal{P}^{d}}wD\left(F\middle\|P\right)+(1-w)D\left(F\middle\|Q\right)$.
A proof sketch can be found in Appendix˜C .
Notice that the test described in Equation˜22 is not symmetric in $P$ and $Q$, while the problem itself is, both in the problem formulation and in the optimal error exponent. Indeed, the result holds when replacing $D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)$ with $D\left(\hat{Q}_{Y^{m}}\middle\|\hat{P}_{X^{n}}\right)$, and or the minimum of the two. Using the minimum test statistic yields a better type I error while achieving the same asymptotic type II exponent.
We now introduce a strong converse for the two-sample problem.
Theorem 5 (Strong Converse).
For a given test $T_{n,m}$, if for a particular $P,Q\in\mathcal{P}^{d}$ with $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$ and $P\neq Q$, we have
$\begin{split}&\liminf_{n,m\to\infty}-\frac{1}{n+m}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in T_{n,m}|(X^{n},Y^{m})\sim(P,Q)\right]\\ &>\left[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)\right].\end{split}$ (26)
Then
$\lim_{n\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\not\in T_{n,m}|X^{n},Y^{m}\sim F^{*}\right]=1.$ (27)
The proof can be found in Appendix˜D
Given $P,Q$, solving for $F^{*}=(f_{k}^{*})_{k}$ yields
$f_{k}^{*}=\frac{p_{k}^{w}q_{k}^{1-w}}{\sum_{k=1}^{d}p_{i}^{w}q_{i}^{1-w}}.$ (28)
Notice that $F^{*}$ is a geometric mixture of the two distributions. It follows that, from [16, Theorem 30], the optimal type II error exponent for the two-sample test is precisely the $w$ Renyi-relative entropy, i.e.,
$wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)=(1-w)D_{w}(P\|Q).$ (29)
In combination Theorem˜4, Theorem˜5, and (29), show that the test in Equation˜22 satisfies
$\begin{split}\lim_{n,m\to\infty}-\frac{1}{n+m}\log\mathbb{P}&\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in A_{n,m}|(X^{n},Y^{m})\sim(P,Q)\right]\\ &=(1-w)D_{w}(P\|Q).\end{split}$ (30)
IV Conclusion
This paper provided a more streamlined proof of the asymptotic optimality of Hoeffding’s likelihood ratio test for the one-sample problem. The test statistic and the proof itself naturally extend to the two-sample test, and we show that a simple test comparing the relative entropy between the two empirical distributions is asymptotically optimal for the two-sample problem. A strong converse was established for the two-sample test: any test whose type II error probability vanishes faster than the optimal type II error exponent necessarily has its type I error approaching one.
Appendix A Preparatory Results
We have the following results that we will need:
Theorem 6 (Pinsker’s Inequality [3]).
$D(P\|Q)\geq\frac{1}{2\ln(2)}\left\|P-Q\right\|_{1}^{2}$ (31)
We also will require a bound on the probability of recovering a particular empirical distribution.
Theorem 7 (Theorem 11.1.4 [3]).
$(n+1)^{-d}2^{-nD\left(\dot{P}\middle\|Q\right)}\leq\mathbb{P}\left[\hat{P}_{X^{n}}=\dot{P}|X^{n}\sim Q\right]\leq 2^{-nD\left(\dot{P}\middle\|Q\right)}$ (32)
Appendix B Proof of Theorem˜3
B-A Type I Error
Under the null hypothesis, we will denote $Q_{X^{n}}$ with $P_{X^{n}}$ to clarify that $P=Q$ here. Now, using Theorem˜1 we have
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ $=\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|P\right)>c_{n}\right]$
$\leq(n+1)^{d}2^{-n\cdot c_{n}}$
Then if $\varepsilon,d$ are fixed, we have
$c_{n}$ $=\frac{d\log(n+1)}{n}+\frac{\log(1/\varepsilon)}{n}$ (33)
$=\Theta\left(\frac{\log(n)}{n}\right)$ (34)
which would cause $\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]\leq\varepsilon$.
B-B Type II Error
If $D\left(P\middle\|Q\right)=+\infty$, then $\exists k\in[d]$ such that $P[k]>0$, but $Q[k]=0$. Here we use the notation $F[k]:=\mathbb{P}_{X\sim F}[X=k]$.
Take $p_{\min}:=\min_{i\in\operatorname{supp}(P)}p_{i}$. Since $c_{n}\searrow 0$ as $n\to\infty$, for $n$ large enough we get
$(2\ln(2))^{1/2}c_{n}^{1/2}<p_{\min}$ (35)
Then if $Q_{X^{n}}\in B_{n}(P,c_{n})$
$\frac{1}{2\ln(2)}\left\|P-Q_{X^{n}}\right\|_{1}^{2}$ $\leq D\left(Q_{X^{n}}\middle\|P\right)\leq c_{n}$ (36)
$\implies\left\|Q_{X^{n}}-P\right\|_{1}\leq(2\ln(2))^{1/2}c_{n}^{1/2}<p_{\min}$ (37)
Therefore, for $n$ large enough, and $Q_{X^{n}}\in B_{n}(P,c_{n})$, we have $P<<Q_{X^{n}}$. But then $Q_{X^{n}}[k]>0$, and since $X^{n}\sim Q$ the probability of seeing such a $Q_{X^{n}}$ is 0 since $Q[k]=0$.
So, if $D\left(P\middle\|Q\right)=+\infty$, and $n$ large enough, we have
$\mathbb{P}\left[D\left(Q_{X^{n}}\middle\|P\right)\leq c_{n}\right]=0$ (38)
Now we prove the result when $D\left(P\middle\|Q\right)<+\infty$. Additionally, without loss of generality we can assume that for all $Q_{X^{n}}\in\mathcal{P}_{n}^{d}$ that $D\left(Q_{X^{n}}\middle\|Q\right)<+\infty$, i.e. $\hat{Q}_{X^{n}}<<Q$.
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ $=\mathbb{P}\left[\hat{Q}_{X^{n}}\in B_{n}(P,c_{n})\right]$ (39)
$\leq(n+1)^{d}2^{-n\ \cdot\min\limits_{Q_{X^{n}}\in B(P,c_{n})}\{D\left(Q_{X^{n}}\middle\|Q\right)\}}$ (40)
Which means we can focus on bounding
$\min\limits_{Q_{X^{n}}\in B(P,c_{n})}\{D\left(Q_{X^{n}}\middle\|Q\right)\}$ (41)
from below.
For $Q_{X^{n}}\in B(P,c_{n})$, we have
$\frac{1}{2\ln(2)}\left\|Q_{X^{n}}-P\right\|_{1}^{2}$ $\leq D(Q_{X^{n}}\|P)\leq c_{n}$ (42)
$\implies\left\|Q_{X^{n}}-P\right\|_{1}$ $\leq(2\ln(2))^{1/2}c_{n}^{1/2}:=r_{n}$ (43)
Since $c_{n}\searrow 0$ we get $r_{n}\searrow 0$. And by our assumptions, we have $D\left(P\middle\|Q\right)<+\infty$, and $D\left(Q_{X^{n}}\middle\|Q\right)<+\infty$. Then invoking Lemma˜1, we have
$\left|D\left(Q_{X^{n}}\middle\|Q\right)-D\left(P\middle\|Q\right)\right|\leq g(Q,r_{n})$ (44)
Now, for $Q$ and $d$ fixed, $g(Q,r_{n})$ is monotonically decreasing in $r_{n}$, and $g(Q,0)=0$. Therefore $\forall\varepsilon>0$, $\exists N>0$ such that $\forall n>N$
$\left|D\left(Q_{X^{n}}\middle\|Q\right)-D\left(P\middle\|Q\right)\right|$ $\leq\varepsilon$ (45)
$\implies D\left(Q_{X^{n}}\middle\|Q\right)$ $\geq D\left(P\middle\|Q\right)-\varepsilon$ (46)
$\implies\min_{Q_{X^{n}}\in B(P,c_{n})}D\left(Q_{X^{n}}\middle\|Q\right)$ $\geq D(P\|Q)-\varepsilon$ (47)
So that for $n$ large enough we have
$\mathbb{P}\left[D\left(Q_{X^{n}}\middle\|P\right)\leq c_{n}\right]$ $\leq(n+1)^{d}2^{-n\left(D(P\|Q)-\varepsilon\right)}$ (48)
$\implies\liminf_{n\to\infty}-\frac{1}{n}\log\beta(A_{n})$ $\geq D\left(P\middle\|Q\right)$ (49)
Appendix C Proof of Theorem˜4
C-A Type I Error
In this section, we will prove that the test described above will asymptotically achieve a Type I error of $\alpha$. By this we mean that given $\alpha\in(0,1)$ for $n,m$ large enough, the test will have Type I error $\leq\alpha$. Under $H_{0}$, we will denote the two empirical distributions by $\hat{P}_{X^{n}}$ and $\hat{P}_{Y^{m}}$, to signify that the underlying distribution for both $X^{n}$ and $Y^{m}$ is $P$.
Now, for any $\delta_{m}>0$, we have
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ $=\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)>c_{m}\right]$ (50)
$=\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)>c_{m}|\hat{P}_{X^{n}}\in B(P,\delta_{m})\right]\mathbb{P}\left[\hat{P}_{X^{n}}\in B(P,\delta_{m})\right]$ (51)
$\ +\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)>c_{m}|\hat{P}_{X^{n}}\in B^{C}(P,\delta_{m})\right]\mathbb{P}\left[\hat{P}_{X^{n}}\in B^{C}(P,\delta_{m})\right]$ (52)
$\leq\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)>c_{m}|\hat{P}_{X^{n}}\in B(P,\delta_{m})\right]+\mathbb{P}\left[\hat{P}_{X^{n}}\in B^{C}(P,\delta_{m})\right]$ (53)
$\leq\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)>c_{m}|\hat{P}_{X^{n}}\in B(P,\delta_{m})\right]+(n+1)^{d}2^{-n\delta_{m}}$ (54)
$\leq\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{n}}\right)>c_{m}|\hat{P}_{X^{n}},\hat{P}_{Y^{m}}\in B(P,\delta_{m})\right]+(n+1)^{d}2^{-n\delta_{m}}+(m+1)^{d}2^{-m\delta_{m}}$ (55)
Since we have $m<n$ by assumption, for $m$ large enough we get
$(n+1)^{d}2^{-n\delta_{n,m}}+(m+1)^{d}2^{-m\delta_{m}}$ $\leq 2(m+1)^{d}2^{-m\delta_{m}}$ (56)
Now, if we take
$\delta_{m}=\frac{1+\log(1/\varepsilon)+d\log(m+1)}{m}=\Theta\left(\frac{\log(m)}{m}\right)$ (57)
which gives us $2(m+1)^{d}2^{-m\delta_{m}}\leq\varepsilon$. Further, takeThe result would hold if we took $c_{m}=\delta_{m}+m^{-\gamma}$ for any $\gamma\in(0,1)$ $c_{m}=\delta_{m}+m^{-1/2}=\Theta(m^{-1/2})$.
Then we can restrict our focus to the event that $\hat{P}_{X^{n}},\hat{P}_{Y^{m}}$ are close to $P$, in the sense that
$D\left(\hat{P}_{X^{n}}\middle\|P\right),D\left(\hat{P}_{Y^{m}}\middle\|P\right)<\delta_{m}.$ (58)
We henceforth analyze $D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)\geq c_{m}$ with this condition, and we will show that for $m$ large enough the probability that both conditions hold is 0.
Using Theorem˜6, we can connect our bounds on the KLD to bounds on the total-variation distance between the two empirical distributions.
$\delta_{m}\geq D\left(\hat{P}_{X^{n}}\middle\|P\right),D\left(\hat{P}_{Y^{m}}\middle\|P\right)$ $\geq\frac{1}{2\ln(2)}\left\|\hat{P}_{X^{n}}-P\right\|_{1}^{2},\left\|\hat{P}_{Y^{m}}-P\right\|_{1}^{2}$ (59)
$\implies\left\|\hat{P}_{X^{n}}-P\right\|_{1},\left\|\hat{P}_{Y^{m}}-P\right\|_{1}$ $\leq\left(2\ln(2)\delta_{m}\right)^{1/2}$ (60)
From [14, Equation 7.34], we have
$D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)$ $\leq\log(e)\cdot\sum_{i=1}^{d}\frac{(\hat{P}_{X^{n}}[i]-\hat{P}_{Y^{m}}[i])^{2}}{\hat{P}_{Y^{m}}[i]}$ (61)
Now, since $\delta_{m}\searrow 0$, $\exists N_{1}>0$ such that $\forall m\geq N_{1}$ we have $\left\|\hat{P}_{Y^{m}}-P\right\|_{1}\leq\frac{p_{\min}}{2}$, so that $\hat{P}_{Y^{m}}[i]\geq\frac{p_{\min}}{2}$. Then assuming $m\geq N_{1}$ we have
$D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)$ $\leq\log(e)\sum_{i=1}^{d}\frac{(\hat{P}_{X^{n}}[i]-\hat{P}_{Y^{m}}[i])^{2}}{\hat{P}_{Y^{m}}[i]}$ (62)
$\leq\frac{2\log(e)}{p_{\min}}\sum_{i=1}^{d}(\hat{P}_{X^{n}}[i]-\hat{P}_{Y^{m}}[i])^{2}$ (63)
$\leq\frac{2\log(e)}{p_{\min}}\left\|\hat{P}_{X^{n}}-\hat{P}_{Y^{m}}\right\|_{1}^{2}$ (64)
$\leq\frac{16}{p_{\min}}\delta_{m}$ (65)
where the last inequality is because
$\left\|\hat{P}_{X^{n}}-\hat{P}_{Y^{m}}\right\|_{1}^{2}$ $\leq 2\left\|\hat{P}_{X^{n}}-P\right\|_{1}^{2}+2\left\|\hat{P}_{Y^{m}}-P\right\|_{1}^{2}$ (66)
$\leq 4\cdot(\delta_{m}2\ln 2)$ (67)
$=8\ln 2\cdot\delta_{m}$ (68)
But since $\delta_{m}=\Theta(\frac{\log m}{m})$ and $c_{m}=\Theta(m^{-1/2})$, $\exists N_{2}>0$ such that
$\frac{16}{p_{\min}}\delta_{m}<c_{m}$ (69)
Take $N_{\max}=\max\{N_{1},N_{2}\}$, then for $m\geq N_{\max}$,
$D\left(\hat{P}_{X^{n}}\middle\|P\right),D\left(\hat{P}_{Y^{m}}\middle\|P\right)$ $\leq\delta_{m}$ (70)
$\implies D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)$ $<c_{m}$ (71)
Therefore, for $m\geq N_{\max}$ we have
$\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{P}_{Y^{m}}\right)\geq c_{m}|\hat{P}_{X^{n}},\hat{P}_{Y^{m}}\in B(P,\delta_{m})\right]=0$ (72)
which gives us our asymptotic type I error bound.
C-B Type II Error
When $\operatorname{supp}(P)\cap\operatorname{supp}(Q)=\emptyset$, then we will always have $D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)=+\infty$, from which we get our first result for the type II error.
Now without loss of generality, we can assume that $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$. Notice that $F^{*}$ is well defined and unique, this was shown previously in [13] but for convenience we show it here in Section˜C-C. Here we will be using $\hat{P}_{X^{n}}$ referring to the empirical distribution of the samples $X^{n}$, and $\hat{Q}_{Y^{m}}$ referring to the empirical distribution of the samples $Y^{m}$.
Let $K_{n,m}=\{(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in\mathcal{P}^{d}_{n}\times\mathcal{P}^{d}_{m}:D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)\leq c_{m}\}$. We have the following argument
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ $=\mathbb{P}\left[D\left(\hat{P}_{X^{n}}\middle\|\hat{Q}_{Y^{m}}\right)<c_{m}\right]$
$=\sum_{\dot{Q}\in\mathcal{P}^{d}_{m}}\ \ \sum_{\dot{P}\in B_{n}(\dot{Q},c_{m})}\mathbb{P}\left[\hat{P}_{X^{n}}=\dot{P},\ \ \hat{Q}_{Y^{m}}=\dot{Q}\right]$
$=\sum_{\dot{Q}\in\mathcal{P}^{d}_{m}}\ \ \sum_{\dot{P}\in B_{n}(\dot{Q},c_{m})}\mathbb{P}\left[\hat{P}_{X^{n}}=\dot{P}\right]\ \ \mathbb{P}\left[\hat{Q}_{Y^{m}}=\dot{Q}\right]$
[…]
$\leq(m+1)^{d}(n+1)^{d}2^{-(n+m)\min_{\begin{subarray}{c}(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in K_{n,m}\end{subarray}}\left[wD\left(\hat{P}_{X^{n}}\middle\|P\right)+(1-w)D\left(\hat{Q}_{Y^{m}}\middle\|Q\right)\right]}$
The third equality is because $X^{n}$ and $Y^{m}$ are independent. The fifth inequality follows from the fact that there are at most $(n+1)^{d}$ and $(m+1)^{d}$ empirical distributions with denominator $n$ or $m$, and from Theorem˜7.
Now, we can focus on bounding
$(P_{n}^{*},Q_{n}^{*})=\operatorname*{arg\,min}_{\begin{subarray}{c}\hat{Q}_{Y^{m}}\in\mathcal{P}^{d}\\ \hat{P}_{X^{n}}\in B(\hat{Q}_{Y^{m}},c_{m})\end{subarray}}\left\{wD\left(\hat{P}_{X^{n}}\middle\|P\right)+(1-w)D\left(\hat{Q}_{Y^{m}}\middle\|Q\right)\right\}$ (73)
Since we assume that $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$, without loss of generality we may assume that $Q_{m}^{*}<<Q$, and $P_{n}^{*}<<Q_{m}^{*}<<Q$. Using the same argument employed in the analysis of the type I error, we must have
$\left\|P_{n}^{*}-Q_{m}^{*}\right\|_{1}\leq(2\ln(2)c_{m})^{1/2}$ (74)
Then invoking Lemma˜1 along with the fact that $c_{m}\searrow 0$ as $m\to\infty$, we get that $\forall\eta>0$, $\exists N>0$ such that
$\left|D\left(P_{n}^{*}\middle\|Q\right)-D\left(Q_{m}^{*}\middle\|Q\right)\right|\leq\eta,\quad\forall m>N$ (75)
Therefore, for $n$ sufficiently large,
$\min\limits_{\begin{subarray}{c}\hat{Q}_{Y^{m}}\in\mathcal{P}^{d}\\ \hat{P}_{X^{n}}\in B_{n}(\hat{Q}_{Y^{m}},c_{m})\end{subarray}}wD\left(\hat{Q}_{Y^{m}}\middle\|Q\right)+(1-w)D\left(\hat{P}_{X^{n}}\middle\|P\right)$ $=wD\left(Q_{m}^{*}\middle\|Q\right)+(1-w)D\left(P_{n}^{*}\middle\|P\right)$ (76)
$\geq wD\left(P_{n}^{*}\middle\|Q\right)-w\eta+(1-w)D\left(P_{n}^{*}\middle\|P\right)$ (77)
$\geq wD\left(F^{*}\middle\|Q\right)+(1-w)D\left(F^{*}\middle\|P\right)-\eta,$ (78)
where the last inequality is because $F^{*}$ is the minimizer of the expression.
So we finally get that $\forall\eta>0$, $\exists N>0$ such that $\forall m\geq N$:
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ $\leq(m+1)^{d}(n+1)^{d}2^{-(n+m)\left[wD\left(F^{*}\middle\|Q\right)+(1-w)D\left(F^{*}\middle\|P\right)-\eta\right]}$ (79)
$\implies\liminf_{n,m\to\infty}-\frac{1}{n+m}\log\beta(A_{n,m})$ $\geq wD\left(F^{*}\middle\|Q\right)+(1-w)D\left(F^{*}\middle\|P\right)$ (80)
C-C Optimal F
We restate the definition of $F^{*}$ here for the sake of clarity,
$F^{*}=\operatorname*{arg\,min}_{F\in\mathcal{P}^{d}}wD\left(F^{*}\middle\|Q\right)+(1-w)D\left(F^{*}\middle\|P\right).$ (81)
Here we are allowed to take the minimum over all probability distributions since $\mathcal{P}^{d}$ is convex and compact, and the function being minimized is convex. We can solve for $F^{*}$ using Lagrangian optimization. Let $F=(f_{1},\dots,f_{d})$ be an optimization variable, and $P=(p_{1},\dots,p_{d})$, $Q=(q_{1},\dots,q_{d})$ be constants (with respect to the optimization). Imposing the constraint $\sum_{i=1}^{d}f_{i}=1$, then our Lagrangian would be
$L(F,\lambda)$ $=wD\left(F\middle\|P\right)+(1-w)D\left(F\middle\|Q\right)-\lambda\left(1-\sum_{i=1}^{d}f_{i}\right)$
$=\sum_{i=1}^{d}f_{i}\left(w\log\left(\frac{f_{i}}{p_{i}}\right)+(1-w)\log\left(\frac{f_{i}}{q_{i}}\right)\right)-\lambda\left(1-\sum_{i=1}^{d}f_{i}\right)$
$=\sum_{i=1}^{d}f_{i}\log\left(\frac{f_{i}}{p_{i}^{w}q_{i}^{1-w}}\right)-\lambda\left(1-\sum_{i=1}^{d}f_{i}\right)$
Now, for each $j\in[d]$, we have
$\frac{\partial}{\partial f_{j}}L(F,\lambda)$ $=\frac{\partial}{\partial f_{j}}\left[f_{j}\log\left(\frac{f_{j}}{p_{j}^{w}q_{j}^{1-w}}\right)\right]+\lambda$
$=\log(f_{j})+\log(e)-\log(p_{j}^{w}q_{j}^{1-w})+\lambda$
Setting this to 0 yields
$f_{j}$ $=p_{j}^{w}q_{j}^{1-w}\cdot 2^{-\left(\log(e)+\lambda\right)}$
And the constraint
$\frac{\partial}{\partial\lambda}L(F,\lambda)$ $=1-\sum_{i=1}^{d}f_{j}$
$=1-2^{-\left(\log(e)+\lambda\right)}\sum_{i=1}^{d}p_{i}^{w}q_{i}^{1-w}$
is satisfied when $2^{-\left(\log(e)+\lambda\right)}=\sum_{i=1}^{d}p_{i}^{w}q_{i}^{1-w}$.
Therefore, our optimal $F^{*}$ has
$f_{j}^{*}=\frac{p_{j}^{w}q_{j}^{1-w}}{\sum_{i=1}^{d}p_{i}^{w}q_{i}^{1-w}}$
Appendix D Proof of Theorem˜5
We first need to establish that the empirical distribution $\hat{P}_{X^{n}}$ is a sufficient statistic for any inference on $P$. We have
$\mathbb{P}\left[X^{n}|X^{n}\sim P\right]=n!\cdot\prod_{i=1}^{d}\frac{(P[i])^{n\cdot\hat{P}_{X^{n}}[i]}}{(n\cdot\hat{P}_{X^{n}}[i])!}.$ (82)
Therefore, by the Neyman-Fisher Factorization Theorem [11], we know that $\hat{P}_{X^{n}}$ is a sufficient statistic for any inference on $P$. And from this we can see that $(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})$ is a sufficient statistic for performing two-sample universal hypothesis testing. This means that we can describe any two-sample universal test in terms of an acceptance region on empirical distributions, $T_{n,m}\subset\mathcal{P}_{d}^{n}\times\mathcal{P}_{d}^{m}$.
Now, we have the following lemma, which says that for a test having type II error rate larger than $wD(F^{*}\|P)+(1-w)D(F^{*}\|Q)$ for some pair of distributions $(P,Q)$, then there must be a gap in $T_{n,m}$ around $(F^{*},F^{*})$ eventually. Let $N_{\eta}(F)=\{G\in\mathcal{P}^{d}:\left\|G-F\right\|_{1}\leq\eta\}$ denote the $\eta$-neighborhood of $F$.
Lemma 2.
Suppose $T_{n,m}$ is a two-sample test satisfying $\exists P,Q\in\mathcal{P}^{d}$ with $P\neq Q$ and $\operatorname{supp}(P)\cap\operatorname{supp}(Q)\neq\emptyset$, so that $\exists\delta>0$ where
$\liminf_{n,m\to\infty}-\frac{1}{n+m}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in T_{n,m}\middle|X^{n}\sim P,Y^{m}\sim Q\right]=wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)+\delta$ (83)
Then $\exists\eta\in(0,1)$, and $M>0$ such that $\forall n,m>M$ we have
$\left(N_{\eta}(F^{*})\times N_{\eta}(F^{*})\right)\cap T_{n,m}=\emptyset.$ (84)
Proof of Lemma˜2.
Suppose for contradiction that $\forall\eta>0$ and $\forall\dot{M}>0$, $\exists n,m>\dot{M}$ with
$\left(N_{\eta}(F^{*})\times N_{\eta}(F^{*})\right)\cap T_{n,m}\neq\emptyset.$ (85)
We will show that this causes a contradiction with the condition on the type II error.
By the statement of the lemma, $\exists M_{\delta}>0$ such that $\forall n,m\geq M_{\delta}$
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in T_{n,m}|X^{n}\sim P,Y^{m}\sim Q\right]\leq 2^{-(n+m)\left[wD(F^{*}\|P)+(1-w)D(F^{*}\|Q)+\frac{\delta}{2}\right]}$
Let $\gamma_{\min}=\min\{p_{\min},q_{\min}\}$ be the minimum of the smallest non-zero probabilities in $P,Q$. Define
$g(x)=x\log(1/\gamma_{\min})+\frac{2\sqrt{2x}}{\ln 2},$ (86)
and take $\eta\in(0,1)$ so that $g(\eta)\leq\frac{\delta}{8}$. Further, let $M_{d}>0$ be such that $\forall m\geq M_{d}$,
$2d\frac{\log n}{n}\leq 2d\frac{\log m}{m}<\frac{\delta}{8}$ (87)
Now, by the conditions we assumed for contradiction, we can take $\dot{M}=M_{\delta}+M_{d}$, and $\exists\dot{n},\dot{m}\geq\dot{M}$ such that $\exists(P_{X^{\dot{n}}},Q_{Y^{\dot{m}}})\in T_{n,m}$ with
$\left\|P_{X^{\dot{n}}}-F^{*}\right\|_{1}$ $\leq\eta$ (88)
$\left\|Q_{Y^{\dot{m}}}-F^{*}\right\|_{1}$ $\leq\eta$ (89)
$P_{X^{\dot{n}}},Q_{Y^{\dot{m}}}$ $<<F$ (90)
Using Lemma˜1 and the condition on $\eta$,
$\left|D\left(P_{X^{\dot{n}}}\middle\|P\right)-D\left(F^{*}\middle\|P\right)\right|\leq g(\eta)$ $\leq\frac{\delta}{8}$ (91)
$\left|D\left(Q_{Y^{\dot{m}}}\middle\|Q\right)-D\left(F^{*}\middle\|Q\right)\right|\leq g(\eta)$ $\leq\frac{\delta}{8}$ (92)
$\implies D\left(P_{X^{\dot{n}}}\middle\|P\right)+D\left(Q_{Y^{\dot{m}}}\middle\|Q\right)$ $\leq D\left(F^{*}\middle\|P\right)+D\left(F^{*}\middle\|Q\right)+\frac{\delta}{4}$ (93)
Since $\dot{n},\dot{m}\geq M_{\delta}$, we also have
$\mathbb{P}\left[(\hat{P}_{X^{\dot{n}}},\hat{Q}_{Y^{\dot{m}}})\in T_{\dot{n},\dot{m}}\middle|X^{\dot{n}}\sim P,Y^{\dot{m}}\sim Q\right]\leq 2^{-(n+m)\left[wD(F^{*}\|P)+(1-w)D(F^{*}\|Q)+\frac{\delta}{2}\right]}$ (94)
But since $(P_{X^{\dot{n}}},Q_{Y^{\dot{m}}})\in T_{\dot{n},\dot{m}}$, we also have
$\mathbb{P}\left[(\hat{P}_{X^{\dot{n}}},\hat{Q}_{Y^{\dot{m}}})\in T_{\dot{n},\dot{m}}|X^{\dot{n}}\sim P,Y^{\dot{m}}\sim Q\right]$ $\geq\mathbb{P}\left[(\hat{P}_{X^{\dot{n}}},\hat{Q}_{Y^{\dot{m}}})=(P_{X^{\dot{n}}},Q_{Y^{\dot{m}}})|X^{\dot{n}}\sim P,Y^{\dot{m}}\sim Q\right]$ (95)
$\overset{(a)}{\geq}2^{-(\dot{n}+\dot{m})\left[wD\left(P_{X^{\dot{n}}}\middle\|P\right)+(1-w)D\left(Q_{Y^{\dot{m}}}\middle\|Q\right)+\frac{2d\log m}{m}\right]}$ (96)
$\overset{(b)}{\geq}2^{-(\dot{n}+\dot{m})\left[wD\left(P_{X^{\dot{n}}}\middle\|P\right)+(1-w)D\left(Q_{Y^{\dot{m}}}\middle\|Q\right)+\frac{\delta}{8}\right]}$ (97)
$\overset{(c)}{\geq}2^{-(\dot{n}+\dot{m})\left[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)+\frac{3\delta}{8}\right]}.$ (98)
Here, $(a)$ comes from Theorem˜7, $(b)$ comes from the fact that $\dot{n},\dot{m}\geq M_{d}$ and Equation˜87, and $(c)$ comes from Equation˜93.
In combination, we have
$2^{-(\dot{n}+\dot{m})\frac{3\delta}{8}}\leq 2^{(\dot{n}+\dot{m})[wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)]}\mathbb{P}\left[(\hat{P}_{X^{\dot{n}}},\hat{Q}_{Y^{\dot{m}}})\in T_{\dot{n},\dot{m}}|X^{\dot{n}}\sim P,Y^{\dot{m}}\sim Q\right]\leq 2^{(\dot{n}+\dot{m})\frac{\delta}{2}}$ (99)
But since $\dot{m}\frac{3\delta}{8}>-\dot{m}\frac{\delta}{2}$, this is a contradiction.
Using Lemma˜2, we can now prove Theorem˜5.
Proof of Theorem˜5.
By the condition in the theorem, $\exists\delta>0$ such that
$\liminf_{n,m\to\infty}-\frac{1}{n+m}\log\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{Q}_{Y^{m}})\in T_{n,m}|X^{n}\sim P,Y^{m}\sim Q\right]=wD\left(F^{*}\middle\|P\right)+(1-w)D\left(F^{*}\middle\|Q\right)+\delta.$ (100)
Then from Lemma˜2, we have $\exists\eta>0$ and $M>0$ such that $\forall n,m\geq M$,
$\left(N_{\eta}(F^{*})\times N_{\eta}(F^{*})\right)\cap T_{n,m}=\emptyset.$ (101)
Now we suppose that $X^{n},Y^{m}\sim F^{*}$, so that $H_{0}$ is true. By the Glivenko-Cantelli theorem [15], we haveThe Glivenko-Cantelli theorem provides the stronger result of almost sure convergence, but for our purposes convergence in probability is sufficient.
$\hat{P}_{X^{n}},\hat{P}_{Y^{m}}\overset{\mathbb{P}}{\to}F^{*},$ (102)
[…]
$\mathbb{P}\left[\left\|\hat{P}_{X^{n}}-F^{*}\right\|_{1}\leq\eta|X^{n}\sim F^{*}\right]$ $\geq 1-\varepsilon$ (103)
$\mathbb{P}\left[\left\|\hat{P}_{Y^{m}}-F^{*}\right\|_{1}\leq\eta|Y^{n}\sim F^{*}\right]$ $\geq 1-\varepsilon$ (104)
$\implies\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in\left(\ N_{\eta}(F^{*})\times N_{\eta}(F^{*})\ \right)|X^{n},Y^{m}\sim F^{*}\right]$ $\geq 1-2\varepsilon$ (105)
Finally, for any $\varepsilon>0$, then $\forall n,m\geq N_{\varepsilon}+M$, we have
$\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in T_{n,m}^{C}|X^{n},Y^{m}\sim F^{*}\right]$ $\geq\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in(N_{\eta}(F^{*})\times N_{\eta}(F^{*}))|X^{n},Y^{m}\sim F^{*}\right]$ (106)
$\geq 1-2\varepsilon,$ (107)
which tells us that
$\lim_{n,m\to\infty}\mathbb{P}\left[(\hat{P}_{X^{n}},\hat{P}_{Y^{m}})\in T_{n,m}^{C}|X^{n},Y^{m}\sim F^{*}\right]=1$ (108)